Assume the following information:

   Percent Complete

Milling DepartmentUnits Materials  Conversion

Beginning work in process inventory 200   40%   30%

Units started into production during March 6,200           

Units completed during the period and transferred to the next department 5,800   100%   100%

Ending work in process inventory 600   70%   60%

Milling DepartmentMaterials Conversion

Cost of beginning work in process inventory$10,000  $15,000  

Costs added during the period 304,000   385,000  

Total cost$314,000  $400,000  

Using the FIFO method, the total cost of the ending work in process inventory with respect to materials is closest to:
Answer #1

Correct answer-----------$43,516

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion Material Conversion
Beginning WIP                   200 40% 30% 60% 70%
Units introduced               6,200
Total units to be accounted for               6,400
Completed and Transferred unit               5,800 0% 0% 100% 100%
Ending WIP                   600 0% 0% 70% 60%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP                   200 60%                         120 70%                              140
From units started/Introduced               5,600 100%                      5,600 100%                          5,600
Total               5,800                      5,720                          5,740
Total Units Material Conversion
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total                   600 70%                         420 60%                              360
Total EUP               6,400                      6,140                          6,100

.

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion TOTAL
Cost incurred in Current Period $ 304,000 $ 385,000 $ 689,000
Total Equivalent Units                                                 6,140                                                         6,100
Cost per Equivalent Units $ 49.51 $ 63.11 $ 112.63
Cost of Beginning WIP $ 10,000 $ 15,000 $ 25,000
Total cost to be accounted for $ 714,000

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred                 5,800.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                        120 $ 49.51 $ 5,941
Conversion                        140 $ 63.11 $ 8,836 $ 14,777
B From units started & completed
Material                     5,600 $ 49.51 $ 277,264
Conversion                     5,600 $ 63.11 $ 353,443 $ 630,706
C=A+B Total $ 645,484
D Cost of Beginning WIP $ 25,000
E=C+D Total Cost of Units transferred $           670,483.90
Cost of Ending WIP                     600.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                        420 $ 49.51 $ 20,795
G Conversion                        360 $ 63.11 $ 22,721
H = F + G Total Cost of Ending WIP $ 43,516
I = E + H Total Cost accounted for $ 714,000
J Rounding Off Differences $ 0
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 714,000